We are experiencing exceptionally hot temperatures over the weekend and these are expected to continue through the week. These high temperatures seem to be impacting some of the electronics that allow us to align and maintain alignment of the primary mirror. The mirror team will continue to improve the actuator zero points, clocking of the segments and gaps between the sensors.
We did not complete the 97 point test last week. This was largely due to teething problems with the new tracker control software. We will continue to work these out this week in preparation for another attempt at this milestone next week. Next week we will also be installing the Video Alignment Telescope and DMI mounts on the corrector test mass to prepare for the alignment of the tracker with the primary mirror. We hope to be able to do both sets of tests next week.
The contractors working on our remote thermal area are making good progress and will continue their preparation for the “vault” cement pour which should take place this Friday. The vault will protect the power and glycol lines as they pass under the service road behind the HET.